OpenStudy (anonymous):

What is the value of coefficient c in this quadratic equation? 5x2 + 6x + 7 = 2(x + 2) 7 4 3 -3 3/5

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

Is it 7?

OpenStudy (amistre64):

no, not 7

OpenStudy (amistre64):

distribute the right side, what do we get?

OpenStudy (anonymous):

2x^2 + 4

OpenStudy (amistre64):

2x + 4 , yes now move everything to the left side and simplify ... what do we get for the constant term?

OpenStudy (anonymous):

yea my fault 2x+4

OpenStudy (anonymous):

5x2 + 6x + 7 = 2(x + 2) There is no "c" character in the above expression.

OpenStudy (anonymous):

4

OpenStudy (amistre64):

7 - 4 is not 4

OpenStudy (amistre64):

and 'c' is the constant term in the quadratic equation: ax^2 + bx + c = 0

OpenStudy (anonymous):

3

OpenStudy (amistre64):

then lets go with 3